# Sequential Quadratic Sum-of-squares Programming for Nonlinear Control Systems

### v1

@@ V-C2 Satellite Attitude Control Synthesis
V-D Reachability Analysis
Our third case deals with the computation of the inner-approximation of the reachable set, using two different dynamical systems, with two and four states respectively. The problem formulation and the derivation can be found in [53, 11]. The primary goal is to synthesize a set in the form $\Omega^{V}_{t,\gamma}=\{x\in\mathbb{R}^{\ell}\mid V(x,t)\leq\gamma\}$, where $V:\mathbb{R}^{\ell}\times\mathbb{R}\rightarrow\mathbb{R}$ is a reachability storage function and $\gamma>0$ bounds the set, together with a time-varying control law $u(t)=\hat{\kappa}(t,x(t))$
We slightly modify the problem by using a similar cost as in the first benchmark, instead of a bisection approach in [53]. We are interested to solve the following SOS problem (cf.[53, 11])
@@ V-D Reachability Analysis
The results are summarized in Table VI. Both algorithms are initialized using the Lyapunov function from the linearized system. For sequential, all multipliers are initialized with simple quadratic polynomials.
TABLE VI: Comparison of sequential SOS (S) and coordinate descent (C) for reachability analysis for different system sizes.
@@
In the first case, sequential needed significantly less iterations. Yet, both approaches needed an equal total time to solve the problem, with CD needing slightly less.
For the 4D example, the sequential SOS outperforms CD. The overall solve time is lower due to the comparable low number of iterations. Sequential was able to solve both problems to a smaller cost. For the second, we were not able to find an optimal sublevel set $\beta$ for coordinate descent. It either turned into infeasibility or was computed to zero by the underlying solver. Thus, for this problem we set $\beta=0.1$ (larger values led to infeasibility), a feasible but potentially conservative, value. The search for the sublevel set in that case could be carried out by an additional bisection which in turn results in a higher computational effort.
V-E Compatible Control Barrier and Control Lyapunov Function Synthesis
The final case addresses the synthesis of compatible Control Barrier Function (CBF) and a Control Lyapunov Function (CLF). Compatible means both conditions are satisfied simultaneously, which is possible by constructing a unified control law that is shared by all CBFs and the CLF [45]. We make use of an adapted version from [45] by also incorporating control constraints and the $\mathcal{K}$ functions. The nonconvex SOS problem reads
[…]
@@ V-E Compatible Control Barrier and Control Lyapunov Function Synthesis
with $\tau=\langle\nabla\hat{V}(x),f(x,\hat{\kappa}(x))\rangle+\gamma_{V}V$ and where $\gamma_{h}$ and $\gamma_{V}$ are extended $\mathcal{K}$ functions of the form a(s)=asa(s)=a\,s with $a>0$. The problem deals with constrained satellite control, with one CBF and one CLF, using the same parameter as from Subsection V-C2. Unlike the previous cases, we concentrate more on the computational aspects of the proposed sequential SOS approach and do not compare to CD. The problem has 1756 constraints and 9404 decision variables. The proposed sequential SOS approach needed 53 iterations to find an optimal solution with optimal cost $J(\hat{V}^{*},\hat{h}^{*},\hat{\kappa}^{*})=0.04$. A detailed breakdown of the computational effort is provided in Table VII. The largest portion is to solve the underlying quadratic subproblem ($67\%$), followed by the filter line-search ($32\%$). The high computational effort of the line-search can be primarily traced back to the constraint violation check using the signed-distance approach ($\approx 70\%$). Other components such as Hessian approximation or function evaluations are in total $\approx 1\%$ and are negligible compared to the SDPs.
TABLE VII: Timing results for the CBF-CLF Satellite Example in seconds.
Total time (Alg. 3) Prob. (3) Alg. 2 Others
77.7 52 24.7 0.9
V-F Discussion and Practical Consideration
The proposed sequential SOS algorithm with its filter line-search allows to directly deal with nonconvex, nonlinear SOS problems, even with non-control affine polynomial system dynamics. Unlike coordinate descent, there is no requirement that the initial guess is a feasible solution. Coordinate descent has in general no convergence guarantees, but local convergence guarantees can be given for the sequential approach under mild assumptions. If the problem is infeasible, confirmed by the feasibility restoration, the user might use these information to adapt the problem, i.e., by changing the polynomial structure or revise the problem formulation. The feasibility restoration adds robustness to the overall algorithm, as demonstrated, for example, by the N-link robot arm example. However, there are no guarantees for the overall algorithm or feasibility restoration to find a feasible point. Whereas a non-feasible initial guess might be advantageous, a bad initial guess might lead to insufficient progress or even divergence since it is a Newtons-based method. The filter line search alleviates the problem, but does not fully resolve it. Strategies for initial guess generation are needed, some of which are outlined next.
@@ V-F Discussion and Practical Consideration
  • Use a simple initial guess (e.g., all polynomials are initialized with basis one) and solve only for a few iterations. Take the result and either use it as an initial guess or use the information to construct an initial guess from it.
@@
  • Remove cost function and try to solve a feasibility problem and/or relax the thresholds for feasibility.
@@
VI Conclusion
This paper demonstrates significant computational savings for nonconvex sum-of-squares problems arising in system analysis and control design are leveraging a filter line-search algorithm. In several real-world inspired case studies, it is demonstrated that the proposed approach outperforms state-of-the-art methods, thereby making a step towards more practical synthesis and analysis tools for system’s control. Besides, of comprehensive benchmarking, this paper is accommodated with an open-source implementation and provides local convergence analysis.
Acknowledgement

### v2

@@ 5.3.2 Spacecraft Attitude Control
5.4 Reachability Analysis
We consider the computation of an inner approximation of the reachable set for two different dynamical systems with two and four states, respectively. The problem formulation and the derivation can be found in Yin2021 ; Cunis2021aut . A state constraint set of the form $\mathcal{X}=\{x\in\mathbb{R}^{n}\mid h(x)\leq 0\}$ is prescribed and the goal is to find a set of the form $\Omega^{V}_{t,\beta}=\{x\in\mathbb{R}^{\ell}\mid V(x,t)\leq\beta\}$, where $V:\mathbb{R}^{\ell}\times\mathbb{R}\rightarrow\mathbb{R}$ is a reachability storage function and $\beta>0$ bounds the set, together with a time-varying control law $u(t)=\hat{\kappa}(t,x(t))$.
We slightly modify the problem by using a similar cost as in the first benchmark instead of the bisection approach in Yin2021 . We solve the nonconvex SOS problem (cf.Yin2021 ; Cunis2021aut )
@@ 5.4 Reachability Analysis
The results are summarized in Table 6. Both algorithms are initialized using the Lyapunov function from the linearized system. For the sequential quadratic SOS approach, all multipliers are initialized with simple quadratic polynomials.
Table 6: Comparison of sequential quadratic SOS (SQ) and coordinate descent (CD) for reachability analysis for different system sizes.
@@
In the 2D example, the sequential quadratic SOS approach needs significantly less iterations, yet both approaches take a similar total time to solve the problem. For the 4D example, the sequential quadratic SOS approach needs significantly less iterations than CD. The overall solve time is lower due to the low number of iterations. The sequential quadratic SOS approach was able to solve both problems to a smaller cost. For the CD, we were not able to find an optimal sublevel set $\beta$ with coordinate descent. Thus, for this example we set $\beta=0.1$ (larger values lead to infeasibility), a feasible but potentially conservative value. The search for the sublevel set could be carried out by an additional bisection which would result in a higher computational effort.
5.5 Compatible Control Barrier and Control Lyapunov Function Synthesis
We consider the synthesis of Control Barrier Function (CBF) and a compatible Control Lyapunov Function (CLF). Here, compatible means both conditions can be satisfied simultaneously, which is shown by constructing a shared control law that for all CBF and the CLF conditions schneeberger_sos_2023 . Adapting from schneeberger_sos_2023 by also incorporating control constraints, we solve the nonconvex SOS problem
$\min_{\begin{subarray}{c}s_{1},s_{2},s_{4}\in\Sigma[x],s_{3}\in\Sigma[x]^{p}\\ \hat{h},\hat{V}\in\mathbb{R}[x],\hat{\kappa}\in\mathbb{R}[x]^{m}\end{subarray}}\;\|g-h\|^{2}_{\mathbb{R}[x]}$ (27a)
@@ 5.5 Compatible Control Barrier and Control Lyapunov Function Synthesis
[…]
Table 7: Timing results for the CBF-CLF Satellite Example in seconds.
Solve time (Alg. 3) Prob. (4) Alg. 2 Others
68.7 46.1 21.9 0.7
5.6 Discussion and Practical Considerations
The sequential quadratic SOS approach with its filter line search allows to directly deal with nonconvex, nonlinear SOS problems, even with noncontrol-affine polynomial system dynamics. Unlike coordinate descent, there is no requirement that the initial guess is a feasible solution and local convergence guarantees can be given for the sequential quadratic SOS approach under mild assumptions. If the problem is infeasible, confirmed by the feasibility restoration, the user might use these information to adapt the problem and change the polynomial structure or revise the problem formulation. The feasibility restoration adds robustness to the overall algorithm, as demonstrated by the $N$-link robot arm example. However, there are no guarantees for the overall algorithm or feasibility restoration to find a feasible point wachterImplementationInteriorpointFilter2006a . Whereas a nonfeasible initial guess might be advantageous, a bad initial guess might lead to insufficient progress or even divergence since it is a Newtons-based method betts2010 . The filter line search alleviates this problem, but does not fully resolve it. Some strategies for initial guess generation are outlined next.
  1. 1.
@@ 5.6 Discussion and Practical Considerations
  2. 2.
    Use a simple initial guess (e.g., all polynomials are initialized with basis one) and solve only for a few iterations. Take the result and either use it as an initial guess or use the information to construct an improved initial guess.
  3. 3.
@@
  4. 4.
    Remove cost function and try to solve a feasibility problem, or relax the thresholds for feasibility.
@@
6 Conclusion
This paper demonstrates significant computational savings for nonconvex sum-of-squares problems arising in system analysis and control design by leveraging a sequential quadratic programming approach with filter line search algorithm. In several real-world-inspired case studies, it is demonstrated that the sequential SOS approach outperforms state-of-the-art methods, thereby making a step towards more practical synthesis and analysis tools for control systems. The results of comprehensive benchmarks are theoretically backed by local superlinear or even quadratic convergence rates under regularity and continuity assumptions. Besides, this paper is accommodated with an open-source implementation.
Acknowledgements.
@@
This research is partially supported by the Ministry of Science, Research and Arts of the state of Baden-Württemberg under funding number MWK32- 7531-49/13/7 for the project DaSO: Data-driven Spacecraft Operations.
The open-source implementation of the proposed filter line search algorithm is available as part of CaΣoS: https://github.com/iFR-OFC/casos.
The data regarding the benchmarks are available in DARUS-5677_2026 .
Declarations
The authors declare that they have no conflict of interest.